# ode]python

### v1

@@ Abstract
Recent advances in machine learning (ML) have accelerated progress in calibrating and operating quantum dot (QD) devices. However, most ML approaches rely on access to large, high-quality labeled datasets for training, benchmarking, and validation, with labels capturing key features in the data. Obtaining such datasets experimentally is challenging due to limited data availability and the labor-intensive nature of labeling. QDFlow is an open-source physics simulator for multi-QD arrays that generates realistic synthetic data with ground-truth labels. QDFlow combines a self-consistent Thomas–Fermi solver, a dynamic capacitance model, and flexible noise modules to produce charge stability diagrams and ray-based data closely resembling experiments. With extensive tunable parameters and customizable noise models, QDFlow supports the creation of large, diverse datasets for ML development, benchmarking, and quantum device research.
@@ 1 Introduction
Among the various quantum computing platforms, quantum dots (QDs) stand out for their scalability potential, compact size, and long coherence times. Operating QD devices, however, remains a formidable challenge, with complexity growing rapidly—often exponentially—as the number of qubits increases. Recent advances in integrating machine learning (ML) with quantum device operation have begun to mitigate these difficulties, offering promising automated control and calibration strategies. For example, ML algorithms have been developed for the fabrication [1, 2], characterization [3, 4], tuning [5, 6, 7, 8, 9, 10, 11, 12], and gate virtualization [13, 14] of QD devices.
@@
Developing robust ML models requires access to large and diverse datasets. Crucially, for supervised ML applications, these datasets must also include metadata that identifies key features, such as the global state (i.e., the number of QD formed), the charge configuration, and the type of transition lines present. Unfortunately, large volumes of high-quality experimental data can be challenging to obtain as companies and research groups often keep such data proprietary [15]. For data that is made publicly available, generating accurate feature labels is a labor-intensive and time-consuming task that produces subjective and, potentially, erroneous labels. Physics-based simulations offer a practical solution: they enable the generation of arbitrarily large datasets while providing direct access to the ground-truth charge states, thereby simplifying the labeling process needed for ML training.
@@
Here, we introduce QDFlow, an open-source Python package for simulating QD systems and generating synthetic data tailored for ML training and applications. The core physics engine in QDFlow employs the Thomas–Fermi approximation to numerically solve for the semi-classical charge density $n(x)$ along a one-dimensional (1D) nanowire. While the current state-of-the-art devices are typically realized by confining charges (electrons or holes) within a two-dimensional (2D) heterostructure, the QDs are ultimately formed within quasi-1D channels within those heterostructures, motivating our choice of a 1D model. In practice, the simulated data produced by QDFlow closely resembles that of linear QD arrays in 2D heterostructures. ML models trained on QDFlow-generated data have been shown to generalize effectively to larger 2D QD arrays [13].
@@
There are several open-source QD device simulators [16, 17, 18], which rely on the constant capacitance model, treating the array of QDs and their associated electrostatic gates as nodes in a network of fixed capacitors. Additionally, Ref. [18] allows the capacitances to vary with respect to the number of charges $n$ by introducing an $n$-dependent correction to the capacitance matrix. In contrast, in QDFlow the capacitance parameters are physics-informed, obtained directly from the self-consistent Thomas–Fermi solution rather than imposed heuristically. All key physical observables—such as current, charge states, and sensor readouts—are derived from a capacitance model constructed based on the computed charge density $n(x)$. This ensures that the capacitances evolve dynamically with gate voltages, yielding a more realistic description of device behavior under tuning. Furthermore, QDFlow allows for modeling regions with low barriers between dots, leading the dots on either side to combine into a single centralized dot. Finally, QDFlow incorporates a flexible noise module, enabling the addition of experimentally relevant effects such as thermal broadening, charge offset drift, and voltage fluctuations. These features make the simulated data qualitatively comparable to experimental measurements while maintaining full access to the ground truth labels required for ML applications. Building on earlier approaches that applied the Thomas–Fermi approximation to model charge densities and stability diagrams [19, 20], QDFlow extends these methods into a flexible, open-source framework that integrates physics-informed capacitance modeling with realistic noise processes tailored for ML applications.
@@
QDFlow has already demonstrated its utility in advancing ML-driven QD research. An earlier, under-development version of the simulator was used to generate the QFlow-lite dataset [21, 20], which enabled the training of several ML models for global state recognition [19, 8, 11]. The dataset also supported the development of a novel classification framework for simple high-dimensional geometrical structures, known as the ray-based classification (RBC) [22]. The expanded dataset, QFlow 2.0: Quantum dot data for machine learning [21], generated using the Thomas-Fermi solver with integrated realistic noise processes, further advanced ML-based approaches to QD tuning. In particular, models trained with data from the QFlow 2.0 dataset have been successfully applied to tasks such as data quality assessment[23], physics-informed RBC and ray-based navigation in 1D QD arrays [12], and the development of a full virtualization stack for 1D and 2D QD arrays [13]. These early successes highlight the value of physics-informed synthetic datasets for accelerating the development of automated control tools for QD systems.
@@
The source code for QDFlow is released under the GNU General Public License in the QDFlow GitHub repository [24]. The library includes full type hints for all classes and functions to ensure clarity, maintainability, and ease of extension.
@@ 2 Physics simulation
QDFlow has three main modules, as depicted in Fig. 1. The core physics engine is contained in the simulation module, which uses a Thomas-Fermi solver to find the stable charge configuration and sensor output of a particular QD device defined by a set of physical parameters. The generate module contains functions to randomize device parameters, as well as to create charge stability diagrams (CSDs) by running the physics simulation many times while changing the gate voltages. This is the module that the user would most often interact with. Finally, the noise module is responsible for adding noise to the final datasets, as well as for applying certain post-processing to the data.
To use QDFlow, a user would first decide whether to use the default configuration already provided or if they want to modify the distribution or range each physics parameter should be randomized over. Then the user would generate one or more sets of randomized device parameters, and for each device, would generate a CSD using the functions in the generate module. QDFlow will then run the physics simulation for each pixel in each diagram, and compile the results into an instance of the CSDOutput dataclass, which it returns to the user. This data, which is in the form of numpy arrays, can immediately be used and plotted, or the user could first add noise to the CSDs to emulate experimental data.
The QDFlow Thomas-Fermi solver was first introduced in Ref. 20, but has since been refined and extended within QDFlow to improve flexibility, physical relevance, and integration with downstream ML workflows. The PhysicsParameters dataclass, which is used to initialize the simulation, specifies over twenty parameters governing the properties of the QD device. These parameters include both material characteristics and device-specific features such as gate geometry and positioning. Importantly, the gate voltages—experimentally relevant control knobs—are explicitly included among the simulation inputs. By sweeping these voltages, QDFlow produces the final outputs: 2D CSDs and 1D rays, directly mirroring the tuning procedures used in real QD experiments.
Once the physics parameters have been specified, an instance of the ThomasFermi class is instantiated. This class serves two main purposes: first, it solves for the charge density function $n(x)$; and second, it uses $n(x)$ to construct a capacitance model and compute physical quantities such as the device’s charge state and sensor response. By repeating this process over a range of gate voltages, QDFlow generates data that can be assembled into CSDs or rays, depending on the application.
@@ 2.1 Nanowire model
QDFlow employs a 1D physics model in which charges are assumed to be confined to a linear nanowire that lies along the $x$-axis. The ends of the nanowire are connected to electron reservoirs, and a bias voltage can be applied between them. Electrostatic gates are positioned at a height $h$ below the xyx\!y-plane, and are modeled as infinite cylindrical conductors with central axis parallel to the $y$-axis, as shown in Fig. 2(a). Gates biased to low potential act as plunger gates, while those biased to high potential act as barrier gates (for positive charge carriers, with the convention reversed for negative carriers).
@@
$V_{\text{gate}}(r)=V_{h}\,\frac{\mathcal{K}_{0}\left(r/\lambda\right)}{\mathcal{K}_{0}\left(h/\lambda\right)}$ (1)
where $V_{h}$ is the potential at a reference distance $r=h$ (chosen as the separation between the gate and the nanowire), $\lambda$ is the screening length, and $\mathcal{K}_{0}(z)$ is the modified Bessel function of the second kind. Specifically, we note that $V_{h}$ is not the voltage of the gate itself; rather, it is the voltage that the nanowire feels due to the gate (in the absence of other gates). This is essentially the voltage of the gate multiplied by the lever arm of the gate.
@@
To make the problem tractable, we adopt the simplifying assumption that the induced charges on each gate are rotationally symmetric. Under this approximation, the induced charges act only as a correction to the central line charge, effectively rescaling $V_{\text{gate}}(r)$ by a constant factor. If we denote the resulting effective potential of gate $i$ as $V^{\prime}_{i}$, the potential $V_{i}$ of gate $i$ is given by:
@@
$A_{ij}=\begin{cases}1&\text{if }i=j\\ V_{\text{gate}}(x_{j}-x_{i})/V_{\text{gate}}(\rho_{j})&\text{otherwise,}\end{cases}$ (3)
where $x_{j}$ is the x-coordinate of the central axis of gate $j$ and $\rho_{j}$ is the radius of the gate $j$. Calculating and inverting the matrix $\mathbf{A}$ allows us to determine the effective potentials $V^{\prime}_{i}$ from the applied gate voltages $V_{i}$. We then obtain $V(x)$ by summing $V_{\text{gate}}\left(\sqrt{(x-x_{i})^{2}+h^{2}}\right)$ over all gates and using the effective potentials $V^{\prime}_{i}$ in place of $V_{h}$ in Eq. (1).
@@ 2.2 Thomas-Fermi solver
$n(x)=\int_{\epsilon^{\prime}(x)}^{\infty}\frac{g_{0}}{1+e^{\beta(\epsilon-\mu)}}d\epsilon=\frac{g_{0}}{\beta}\operatorname{sp}\left[\beta(\mu-\epsilon^{\prime}(x))\right]$ (4)
$\epsilon^{\prime}(x)=qV(x)+\int_{\mathds{R}}K(x,x^{\prime})n(x^{\prime})dx^{\prime}$ (5)
The physical parameters $\mu$, $g_{0}$, and $\beta$ in Eq. (4) indicate the Fermi level, the density of states in the conduction band (which is constant in 2D), and the inverse temperature, respectively, and $\operatorname{sp}(z)=\ln(1+e^{z})$ is the softplus function. Parameter $q$ in Eq. (5) controls the sign of the charge carriers, with 1-1 indicating electrons and +1+1 indicating holes, while $K(x,x^{\prime})$ gives the strength of the Coulomb interaction between points $x$ and $x^{\prime}$, and is defined as follows:
$K(x,x^{\prime})=\frac{K_{0}}{\sqrt{(x-x^{\prime})^{2}+\sigma^{2}}}$ (6)
where $K_{0}$ defines the energy scale of the interaction and $\sigma$ is a softening parameter added to prevent a numerical singularity at $x=x^{\prime}$, which occurs due to the 1D model breaking down at scales less than the radius of the nanowire. The value of $\sigma$ can be chosen to be $3\pi r/8$, where $r$ is the nanowire radius, to maintain consistency with the potential energy of two uniformly charged disks as the spacing between them approaches zero [25], or alternatively, a custom interaction $K(x,x^{\prime})$ can be provided.
The Coulomb integral in Eq. (5) is formally taken over the entire nanowire. Because $K(x,x^{\prime})$ scales for large $x^{\prime}$ as $1/|x^{\prime}|$, this introduces concerns that the integral might diverge. At the same time, the integrand is weighted by $n(x^{\prime})$, which becomes exponentially small for $V(x^{\prime})-\mu\gg\beta^{-1}$. This condition is satisfied at the external barriers, see Fig. 2(b), provided that the external barrier voltages are sufficiently high. In addition, it is assumed that at distances far away from the nanowire, the system is connected to an electron reservoir where $n(x^{\prime})$ is large. However, the Coulomb interaction in semiconductors tends to include a screening term that suppresses contributions past a certain range. Thus, in practice, it is sufficient to evaluate the integral between the two external barrier gates.
@@
$n(x)=\frac{g_{0}}{\beta}\operatorname{sp}\left[\beta(\mu-qV(x)-\mathbf{K}n(x))\right].$ (8)
@@
The convergence can be problematic for certain parameter regimes. For example, if we define $\Delta(x)$ to be the difference between an initial guess $n_{0}(x)$ and the true value $n(x)$:
$n_{0}(x)=n(x)+\Delta(x),$ (9)
@@
$n_{1}(x)=\frac{g_{0}}{\beta}\operatorname{sp}\left[\beta(\mu-qV(x)-\mathbf{K}n(x)-\mathbf{K}\Delta(x))\right]$ (10)
We now use the approximation $\operatorname{sp}(z)\approx z$, which is valid for $z\gg 1$. Although this assumption does not always hold (particularly for small $\beta$), it is useful for analyzing certain convergence issues that may arise. Under this approximation, Eq. (10) simplifies to:
$n_{1}(x)\approx n(x)-g_{0}\mathbf{K}\Delta(x)$ (11)
If all eigenvalues of $g_{0}\mathbf{K}$ are smaller than 1, the error term -$\text{-}g_{0}\mathbf{K}\Delta(x)$ will be smaller in magnitude than the initial error $\Delta(x)$, and successive iterations will therefore converge to $n(x)$. Conversely, if $g_{0}\mathbf{K}$ possesses eigenvalues greater than 1, the iterative scheme will generally diverge. Physically, this divergence corresponds to strong coupling between charges, a regime that is well known to cause convergence difficulties in condensed matter systems [26]. Fortunately, this issue can be partially mitigated by solving Eq. (9) for $\Delta(x)$, substituting the result into Eq. (11), and solving for $n(x)$, yielding the following expression:

### v2

@@ Abstract
Recent advances in machine learning (ML) have accelerated progress in calibrating and operating quantum dot (QD) devices. However, most ML approaches rely on access to large, representative datasets designed to capture the full spectrum of data quality encountered in practice, with both high- and low-quality data for training, benchmarking, and validation, with labels capturing key features of the device state. Collating such datasets experimentally is challenging due to limited data availability, slow measurement bandwidths, and the labor-intensive nature of labeling. QDFlow is an open-source physics simulator for multi-QD arrays that generates realistic synthetic data with ground-truth labels. QDFlow combines a self-consistent Thomas-Fermi solver, a dynamic capacitance model, and flexible noise modules to simulate charge stability diagrams and ray-based data closely resembling experiments. With an extensive set of parameters that can be varied and customizable noise models, QDFlow supports the creation of large, diverse datasets for ML development, benchmarking, and quantum device research.
@@ 1 Introduction
Among the various quantum computing platforms, quantum dots (QDs) stand out for their scalability potential, compact size, and long coherence times [Burkard21-SSQ]. Operating QD devices, however, remains a formidable challenge, with complexity growing rapidly—often exponentially—as the number of qubits increases. Recent advances in integrating machine learning (ML) with quantum device operation have begun to mitigate these difficulties, offering promising automated control and calibration strategies. For example, ML algorithms have been developed for the fabrication [1, 2], characterization [3, 4], tuning [5, 6, 7, 8, 9, 10, 11, 12], and gate virtualization [13, 14] of QD devices.
@@
Developing robust ML models requires access to large and diverse datasets representative of the multi-dimensional parameter space typical of QD devices. Crucially, for supervised ML applications, these datasets must also include metadata that identifies key features, such as the global state (i.e., the number of QDs formed), the charge configuration, and the type of transition lines present. Unfortunately, large volumes of high-quality experimental data can be challenging to obtain as companies and research groups often keep such data proprietary [15]. Limited measurement bandwidth in real-world experiments also constrains the efficient exploration of the entire high-dimensional parameter space in a reasonable time. Generating accurate feature labels for publicly available data is a labor-intensive and time-consuming task that can produce subjective and potentially erroneous labels. Physics-based simulations offer a practical solution: they enable the generation of arbitrarily large datasets while providing direct access to the ground-truth charge states, thereby simplifying the labeling process needed for ML training.
@@
Here, we introduce QDFlow, an open-source Python package for simulating QD systems and generating synthetic data tailored for ML training and applications. The core physics engine in QDFlow employs the Thomas-Fermi approximation to numerically solve for the semi-classical charge density $n(x)$ along a one-dimensional (1D) nanowire. While the current state-of-the-art devices are typically realized by confining charges (electrons or holes) within a two-dimensional (2D) heterostructure, the QDs are ultimately formed within quasi-1D channels within those heterostructures, motivating our choice of a 1D model. In practice, the simulated data produced by QDFlow closely resembles that of linear QD arrays in 2D heterostructures. ML models trained on QDFlow-generated data have been shown to generalize effectively to larger 2D QD arrays [13].
@@
There are several open-source QD device simulators [16, 17, 18], which rely on the constant capacitance model, treating the array of QDs and their associated electrostatic gates as nodes in a network of fixed capacitors. Additionally, Ref. [18] allows the capacitances to vary with respect to the number of charges $n$ by introducing an $n$-dependent correction to the capacitance matrix. In contrast, in QDFlow the capacitance parameters are physics-informed, obtained directly from the self-consistent Thomas-Fermi solution rather than imposed heuristically. All key physical observables—such as current, charge states, and sensor readouts—are derived from a capacitance model constructed based on the computed charge density $n(x)$. This ensures that the capacitances evolve dynamically with gate voltages, yielding a more realistic description of device behavior. Furthermore, QDFlow allows for modeling regions with low barriers between dots, leading the dots on either side to combine into a single centralized dot. Finally, QDFlow incorporates a flexible noise module, enabling the addition of experimentally relevant effects such as thermal broadening, charge offset drift, and voltage fluctuations. These features make the simulated data qualitatively comparable to experimental measurements while maintaining full access to the ground truth labels required for ML applications. Building on QFlow—a legacy implementation of the QD simulator that applied the Thomas-Fermi approximation to model charge densities and stability diagrams [19, 20]QDFlow extends these methods into a flexible, open-source framework that integrates physics-informed capacitance modeling with realistic noise processes tailored for ML applications.
@@
QDFlow has already demonstrated its utility in advancing ML-driven QD research. The legacy version of the simulator was used to generate the QFlow-lite dataset [21, 20], which enabled the training of several ML models for global state recognition [19, 8, 11]. The dataset also supported the development of a novel classification framework for simple high-dimensional geometrical structures, known as the ray-based classification (RBC) framework [22]. The expanded dataset, QFlow 2.0: Quantum dot data for machine learning [21], generated using the Thomas-Fermi solver with integrated realistic noise processes, further advanced ML-based approaches to QD tuning. In particular, models trained with data from the QFlow 2.0 dataset have been successfully applied to tasks such as data quality assessment[23], physics-informed RBC and ray-based navigation in 1D QD arrays [12], and the development of a full virtualization stack for 1D and 2D QD arrays [13]. These early successes highlight the value of physics-informed synthetic datasets for accelerating the development of automated control tools for QD systems.
@@
QDFlow is publicly available for download from the Python Package Index, with the source code released under the GNU General Public License and comprehensive documentation and tutorials to support new users included in the QDFlow GitHub repository [24]. The library includes full type hints for all classes and functions to ensure clarity, maintainability, and ease of extension.
@@ 2 Physics simulation
QDFlow has three main modules: simulation, noise, and generate, and one utility module, distribution, as depicted in Fig. 1. The simulation and noise modules are part of the physics package. The core physics-based engine of the simulator is contained in the simulation module. It uses a Thomas-Fermi solver to find the stable charge configuration and sensor output of a particular QD device defined by a set of physical parameters. The QDFlow Thomas-Fermi solver was first introduced in Ref. 20, but has since been refined and extended within QDFlow to improve flexibility, physical relevance, and integration with downstream ML workflows. The PhysicsParameters dataclass, which is used to initialize the simulation, specifies over twenty parameters governing the properties of the QD device. These parameters include both material characteristics and device-specific features such as gate geometry and positioning. Importantly, the gate voltages—experimentally relevant control knobs—are explicitly included among the simulation inputs. By sweeping these voltages, QDFlow produces the final outputs: 2D CSDs and 1D rays, directly mirroring the tuning procedures used in real QD experiments.
The noise module is responsible for adding noise to the final datasets, as well as for applying certain post-processing to the data. The generate module contains high-level functions to assist in generating datasets. It is the module that the user would most often interact with. Finally, the distribution module, contained within the util package, contains classes defining random variable distributions.
To generate data with QDFlow, the user first chooses whether to run the default configuration or adjust the distributions and ranges over which physics parameters are randomized. Next, they create one or more sets of randomized device parameters, and for each device, generate a CSD using the functions in the generate module. Once the physics parameters have been specified, an instance of the ThomasFermi class is instantiated. This class serves two main purposes: first, it solves for the charge density function $n(x)$; and second, it uses $n(x)$ to construct a capacitance model and compute physical quantities such as the device’s charge state and sensor response. QDFlow then runs the physics simulation for every pixel in each diagram and compiles the results into a CSDOutput dataclass, which is returned to the user. By repeating this process over a range of gate voltages, QDFlow generates data that can be assembled into CSDs or rays, depending on the application. The output, stored as NumPy arrays, can be directly analyzed and plotted, or optionally augmented with noise to emulate experimental data.
In the following sections, we provide a more detailed account of the nanowire model physics underlying the simulation. We then explain how the Thomas-Fermi approximation is applied to construct the capacitance model that drives the CSD simulation.
@@ 2.1 Nanowire model
QDFlow employs a 1D physics model in which charges are assumed to be confined to a linear nanowire that lies along the $x$-axis. The ends of the nanowire are connected to electron reservoirs, and a bias voltage can be applied between them. Electrostatic gates are positioned at a height $h$ below the xyx\!y-plane, and are modeled as infinite cylindrical conductors with central axis parallel to the $y$-axis, as shown in Fig. 2(a). The arrangement makes our nanowire model a hybrid between a true nanowire device and other QD device architectures. Gates biased to low potential act as plunger gates, while those biased to high potential act as barrier gates (for positive charge carriers, with the convention reversed for negative carriers).
@@
[…]
where $V_{h}$ is the potential at a reference distance $r=h$ (chosen as the separation between the gate and the nanowire), is the screening length, and $\mathcal{K}_{0}(z)$ is the modified Bessel function of the second kind. Specifically, we note that $V_{h}$ is not the voltage of the gate itself; rather, it is the voltage that the nanowire feels due to the gate (in the absence of other gates). This is essentially the voltage of the gate multiplied by the lever arm of the gate.
@@
To make the problem tractable, we adopt the simplifying assumption that the induced charges on each gate are rotationally symmetric about the axis of the gate. Under this approximation, the induced charges act as a line charge that lies along the central axis of the gate. Because the gate is rotationally symmetric, the potential $V_{\text{gate}}(r)$ also acts as a single line charge at the center of the gate, and thus the induced charges effectively rescale $V_{\text{gate}}(r)$ by a constant factor. Let $V^{\prime}_{i}$ be the rescaled value of $V_{h}$ for gate $i$ after including the effects of the induced charges on gate $i$, and let $V_{i}$ be the value of $V_{h}$ necessary for $V_{\text{gate}}(r)$ to give the actual potential of gate $i$. Then by the superposition principle, $V_{i}$ can be determined by adding the potential contributions from each of the gates:
@@
where $A_{ij}$ denotes the ratio between the contribution from gate $j$ to the potential at gate $i$ and the effective potential $V^{\prime}_{j}$. In the absence of other gates, $V_{i}=V^{\prime}_{i}$, so $A_{ii}$ is simply 1. The contribution from one gate to another is given by $V_{\text{gate}}(r)$, where $r$ is the distance between gates. Thus, $A_{ij}$ can be expressed as follows:
$A_{ij}=\begin{cases}1&\text{if }i=j\\ V_{\text{gate}}(x_{j}-x_{i})/V_{\text{gate}}({}_{j})&\text{otherwise,}\end{cases}$ (3)
where $x_{j}$ is the x-coordinate of the central axis of gate $j$ and j is the radius of the gate $j$. Calculating and inverting the matrix $\mathbf{A}$ allows us to determine the effective potentials $V^{\prime}_{i}$ from the applied gate voltages $V_{i}$. We then obtain $V(x)$ by summing $V_{\text{gate}}\left(\sqrt{(x-x_{i})^{2}+h^{2}}\right)$ over all gates and using the effective potentials $V^{\prime}_{i}$ in place of $V_{h}$ in Eq. (1).
@@ 2.2 Thomas-Fermi solver
[…]
${}^{\prime}(x)=qV(x)+\int_{\mathds{R}}K(x,x^{\prime})n(x^{\prime})dx^{\prime}$ (5)
The physical parameters , $g_{0}$, and in Eq. (4) indicate the Fermi level, the density of states in the conduction band (which is constant in 2D), and the inverse temperature, respectively, and $\operatorname{sp}(z)=\ln(1+e^{z})$ is the softplus function. Parameter $q$ in Eq. (5) controls the sign of the charge carriers, with 1-1 indicating electrons and +1+1 indicating holes, while $K(x,x^{\prime})$ gives the strength of the Coulomb interaction between points $x$ and $x^{\prime}$, and is defined as follows:
$K(x,x^{\prime})=\frac{K_{0}}{\sqrt{(x-x^{\prime})^{2}+{}^{2}}}$ (6)
[…]
@@
[…]
@@
[…]
@@
[…]
We now use the approximation $\operatorname{sp}(z)\approx z$, which is valid for $z\gg 1$. Although this assumption does not always hold (particularly for small ), it is useful for analyzing certain convergence issues that may arise. Under this approximation, Eq. (10) simplifies to:
[…]